### v1

# Quantum Energy Teleportation under Equilibrium and Nonequilibrium Environments

Abstract
Quantum energy teleportation (QET), implemented via local operations and classical communication, enables carrier-free energy transfer by exploiting quantum resources. While QET has been extensively studied theoretically and validated experimentally in various quantum platforms, enhancing energy output for mixed initial states, as the system inevitably interacts with environments, remains a significant challenge. In this work, we study QET performance in a two-qubit system coupled to equilibrium or nonequilibrium reservoirs. We derive an analytical expression for the energy output in terms of the system Hamiltonian eigenstates, enabling analysis of energy output for mixed states. Using the Redfield master equation, we systematically examine the effects of qubit detuning, nonequilibrium temperature difference, and nonequilibrium chemical potential difference on the energy output. We find that the energy output for mixed states often follows that of the eigenstate with the highest population, and that nonequilibrium environments can enhance the energy output in certain parameter regimes.
I Introduction
Quantum teleportation (QT) is a well-known protocol that transmits the information of quantum states to remote locations using quantum entanglement together with local operations and classical communication (LOCC) [1, 2, 3]. Later, Hotta introduced a novel protocol called quantum energy teleportation (QET), which enables energy transmission via entanglement and LOCC [4]. Theoretically, QET can be realized in various physical systems, including spin chains [5, 6, 7], cold trapped ions [6], harmonic chains [8], and quantum fields [4, 9, 10, 11, 12]. QET has also been studied in holographic conformal field theory [13]. Recently, QET was experimentally demonstrated in the laboratory and on a quantum chip [14, 15]. Although QT and QET both rely on quantum correlations, their goals differ: QT transmits quantum state information, whereas QET aims to extract energy from a local subsystem rather than to restore the state.
In the original QET protocol [4], the sender (Alice) and the receiver (Bob) share the ground state, so Bob cannot extract energy from his subsystem by any local unitary operation alone. Alice’s measurement, however, increases the energy of her subsystem, and Bob’s subsystem collapses into a state that is no longer of minimal energy due to entanglement, enabling him to extract energy. The extracted energy is nevertheless less than the energy Alice injected, ensuring energy conservation. Alice must transmit her measurement outcome to Bob via classical communication before he can perform the extraction. Because there is no physical carrier transferring energy between Alice and Bob, the effective speed of energy transfer is limited only by the speed of classical communication.
If Bob’s operations may include arbitrary local operations, he can extract more energy; this is referred to as strong QET [16]. Strong QET consistently yields greater extracted energy than standard QET, and they coincide only when Bob’s subsystem is left in a pure state after Alice’s measurement. Recently, Kazuki Ikeda proposed extending QET concepts beyond energy to arbitrary observables [17]. To illustrate this idea, he studied a (1+1)-dimensional Dirac system and used feedback control based on fermion chirality to activate electric current and charge, and he derived a rigorous upper bound on the teleported quantity. In conventional QET, the upper bound on energy output is severely constrained by distance; however, using squeezed vacuum states with local vacuum regions between the two parties can overcome this limitation [18]. In addition, a hyperbolic quantum network can realize long-range QET by transmitting local quantum information via quantum teleportation and performing conditional operations on that information [19].
Energy teleportation necessarily requires quantum resources, but the specific resources relevant to QET depend on the setting and without an universal consensus. In the minimal QET model, Lin et al. found that initial-state entanglement and coherence show no clear relationship with the extractable energy, although they correlate positively with the energy-output efficiency [20]. Moreover, the change in system entropy during the measurement process sets a lower bound on the transferable energy [7]. For thermal states, QET is enabled by thermal discord [21]. However, in some cases quantum discord is not the resource for QET, as shown for a three-spin Ising chain in a Gibbs state [22]. Note that the total amount of transmitted energy and information is constrained by entanglement [23].
In practice, QET protocols inevitably involve environmental interactions. Since Alice and Bob are located separately, it is essential to account for the effects of distinct local environments on QET. In the standard QET model the system is assumed to be in the ground state, which yields both a low total energy transfer and low efficiency [24]. However, when the system is in a mixed state, the presence of excited populations need not be detrimental to energy extraction. Here we consider QET in a two-qubit model where each qubit interacts only with its own environment, and we investigate how equilibrium and nonequilibrium reservoirs can be exploited to improve QET performance.
Notably, previous studies have shown that nonequilibrium environments can enhance various of quantum correlations, including the quantum entanglement [25, 26, 27, 28], quantum discord [29, 30, 31], quantum steering [32], Bell nonlocality [33], and temporal correlations [34, 35]. We consider QET under steady-state conditions, where, after completing a total protocol, the environment has “cooled” the system back to its initial state. Nonequilibrium steady states exhibit properties distinct from equilibrium cases [36, 37, 38, 39]. We apply the Bloch–Redfield master equation to describe the nonequilibrium two-qubit model, which enables us to simulate changes in energy output due to temperature or chemical potential differences between the baths [40, 41, 42, 43, 44, 45, 46]. Compared with the Lindblad master equation, the Bloch–Redfield equation, without the secular approximation, provides a more accurate description of nonequilibrium steady states [47, 48, 49, 50, 51, 52, 53]. The limitations of the Redfield equation regarding density-matrix positivity and methods to mitigate this issue are discussed in [43, 46, 54, 55]. Additionally, we consider the effect of detuning between the system energy levels, which can enhance nonequilibrium effects.
We find that a temperature difference in bosonic reservoirs consistently suppresses QET, whereas in fermionic reservoirs a temperature difference can enhance QET. The chemical potential difference has a strong effect: when the average chemical potential is extreme (either much smaller or much larger than the system energy levels), QET is reduced; conversely, when the chemical potential is comparable to the system energy levels, QET can be enhanced within a certain range. For a system in a low-excitation state, increasing Alice’s energy level can improve the energy output, while for a system in a high-excitation state, increasing Bob’s energy level can likewise raise the energy output.
The paper is organized as follows. In Sec. II we introduce the standard QET protocol and analyze the energy output when QET is performed on each eigenstate of the Hamiltonian. We also review the Redfield master equation used in our study. QET under equilibrium and nonequilibrium environments is analyzed in Secs. III and IV, respectively. Finally, in Sec. V we summarize our findings. For simplicity, we set $\hbar=k_{B}=1$ in the following sections.
II Energy Teleportation and Redfield Equation
In this section, we first review the protocol of QET in Sec. II.1. Next, we analyze the energy output of the initial mixed state with an ”X” structure in Sec. II.2. Finally, we establish the model for our study in Sec. II.3, namely two qubits coupled to nonequilibrium environments.
II.1 Two-qubit Model of Energy Teleportation
The minimal QET model, known as the two-particle Hotta model [56], considers interacting Heisenberg spin-1/2 particle pair as qubits A and B, possessed by Alice and Bob, respectively. The Hamiltonian of this model in the standard QET protocol is designed with zero ground energy, and the protocol is performed in ground state. For the general QET protocol, the initial state is not the entangled ground state but rather an arbitrary quantum state [16]. The Hamiltonian of the system is set as
$H_{AB}=H_{A}+H_{B}+V=\varepsilon_{A}\sigma^{z}_{A}+\varepsilon_{B}\sigma^{z}_{B}+2\kappa\sigma^{x}_{A}\sigma^{x}_{B},$ (1)
where $\varepsilon_{A,B}$ are the energy levels; $\kappa$ is interaction strength between the qubit $A$ and $B$; $\sigma^{z}_{A,B}$ and $\sigma^{x}_{A,B}$ are the Pauli operators of the qubits $A$ and $B$. In the original QET model, the energy level $\varepsilon_{A}$ equal to $\varepsilon_{B}$. We relax this constraint in our study, and consider detuning of the energy levels as an asymmetrical condition of system. Correspondingly the energy of the ground state is not necessarily zero [16]. The eigenvalues of the Hamiltonian in Eq. (1) are
$E_{1}$ $=-\sqrt{\Omega^{2}+4\kappa^{2}},$
$E_{2}$ $=-\sqrt{\Delta^{2}+4\kappa^{2}},$
$E_{3}$ $=\sqrt{\Delta^{2}+4\kappa^{2}},$
$E_{4}$ $=\sqrt{\Omega^{2}+4\kappa^{2}},$ (2)
and the corresponding eigenstates are
$|E_{1}\rangle$ $=-\sin\phi_{1}|11\rangle+\cos\phi_{1}|00\rangle,$
$|E_{2}\rangle$ $=-\sin\phi_{2}|10\rangle+\cos\phi_{2}|01\rangle,$
$|E_{3}\rangle$ $=\cos\phi_{2}|10\rangle+\sin\phi_{2}|01\rangle,$
$|E_{4}\rangle$ $=\cos\phi_{1}|11\rangle+\sin\phi_{1}|00\rangle,$ (3)
where $\Omega=\varepsilon_{A}+\varepsilon_{B}$ and $\Delta=\varepsilon_{A}-\varepsilon_{B}$. The angles $\phi_{1}$ and $\phi_{2}$ are given by
$\phi_{1}$ $=\arctan\left(\frac{2\kappa}{\Omega+\sqrt{\Omega^{2}+4\kappa^{2}}}\right),$
$\phi_{2}$ $=\arctan\left(\frac{2\kappa}{\Delta+\sqrt{\Delta^{2}+4\kappa^{2}}}\right).$ (4)
The QET protocol consists of three steps [56]: (i) First Alice performs projective measurements
$P_{A}(u)=\frac{1}{2}(I+u\sigma^{x}_{A}),$ (5)
on her qubit A and obtains the results $u\in{\{\pm 1\}}$; (ii) Then Alice communicates the measurement result $u$ to Bob via a classical channel; (iii) Bob performs a local unitary operation $U_{B}(u)$ based on the value of $u$. The operation $U_{B}(u)$ is given by
$U_{B}(u)=I\cos\theta-iu\sigma^{y}_{B}\sin\theta,$ (6)
where $\theta$ is an adjustable real number.
We generalize the original QET scenario from an initial pure state to a mixed state, denoted as $\rho_{AB}$. The initial state has the energy
$E_{0}(\rho_{AB})=\mathrm{Tr}(H_{AB}\rho_{AB}).$ (7)
After Alice performs projective measurements $P_{A}(u)$, the expected energy of the system is given by
$E_{A}(\rho_{AB})=\sum_{u=\pm 1}\mathrm{Tr}\left(H_{AB}P_{A}(u)\rho_{AB}P_{A}^{\dagger}(u)\right).$ (8)
The measurements $P_{A}(u)$ only affect the energy of subsystem $A$, while the energy of subsystem $B$ remains unchanged, due to
$[P_{A}(u),H_{B}]=[P_{A}(u),V]=0.$ (9)
After Alice sends the measurement result of $u$ to Bob, then Bob performs $U_{B}(u)$ on his qubit. Then the energy of system becomes
$E_{B}(\rho_{AB})=\sum_{u=\pm 1}\mathrm{Tr}\left(H_{AB}U_{B}(u)P_{A}(u)\rho_{AB}P_{A}^{\dagger}(u)U_{B}^{\dagger}(u)\right).$ (10)
The energy difference $E_{\text{out}}=E_{A}-E_{B}$ represents as the energy output to $B$ with the help of $A$.
Suppose that the mixed initial state $\rho_{AB}$ is a classical mixture of four eigenstates, such as the thermal state. Before analyze the QET on the mixture $\rho_{AB}$, we first calculate the energy output $E_{\text{out}}$ of four eigenstates
$E_{\text{out}}(|E_{1}\rangle)=-E_{\text{out}}(|E_{4}\rangle)\\ =\frac{1}{\sqrt{\Omega^{2}+4\kappa^{2}}}\left(2\varepsilon_{A}\kappa\sin 2\theta-(\varepsilon_{B}\Omega+4\kappa^{2})(1-\cos 2\theta)\right),$ (11)
$E_{\text{out}}(|E_{2}\rangle)=-E_{\text{out}}(|E_{3}\rangle)\\ =\frac{1}{\sqrt{\Delta^{2}+4\kappa^{2}}}\left(-2\varepsilon_{A}\kappa\sin 2\theta+(\varepsilon_{B}\Delta-4\kappa^{2})(1-\cos 2\theta)\right).$ (12)
Clearly, as the parameter $\theta$ varies (from the correction $U_{B}(u)$), the energy output also changes. However, there is no single optimal parameter $\theta$ that can maximize all values of $E_{\text{out}}$ simultaneously, as illustrated in Fig. 1.
As the original protocol designed for the ground state, its application to excited states presents certain inconsistencies. For instance, when the $\theta$ of QET protocol from ground state is applied to the first excited state, the energy output $E_{\text{out}}$ is found to be less than that in the ground state scenario. Furthermore, the calculated values of $E_{\text{out}}$ for the third and highest excited states are negative, as shown in Fig. 1. This is an intriguing phenomenon, as it indicates that higher energy require a modification of the protocol. We discovered that by adjusting the parameter $\theta$ in the protocol, the calculated energy output values for the third and highest excited states can become positive; however, in this case, the energy output for the ground and first excited states turn negative.
It is evident that $E_{\text{out}}(|E_{1}\rangle)(E_{\text{out}}(|E_{2}\rangle))$ and $E_{\text{out}}(|E_{4}\rangle)(E_{\text{out}}(|E_{3}\rangle))$ exhibit opposite behaviors from Eq. (12) and Fig. 1. This implies that when we select the parameter $\theta$ to maximize $E_{\text{out}}(|E_{1(2)}\rangle)$, the corresponding value of $E_{\text{out}}(|E_{4(3)}\rangle)$ is minimized. Therefore, in the case of mixed states, the maximum energy output is determined by the density matrix resulting from the superposition of the four eigenstates. However, when a specific state dominates (i.e., its proportion is high), the behavior of $E_{\text{out}}$ closely resembles that of this state. This enables a qualitative analysis of $E_{\text{out}}$ under specific conditions.
Figure 1: Energy output of four eigenstates of $H_{AB}$ (1). The parameters are set as $\kappa=1$ and $\varepsilon_{A}=\varepsilon_{B}=2$.
II.2 Energy teleportation with $X$ state
Apparently the energy output originates from the correlation between A and B. However, the specific quantum resources underpinning energy teleportation still lack a comprehensive explanation. For the ground state, the efficiency of energy transfer is closely related to coherence and concurrence [20]. However, in the case of mixed states, it remains unclear which specific quantum resources fully determine energy transfer. While the presence of quantum resources allows for greater energy output, nonetheless, the total amount of energy extracted and the efficiency of extraction do not always vary monotonically with respect to any specific quantum resource. In more extreme cases, it is possible to extract energy even when performing QET on a direct product state [57]. Although a unique relationship between energy output and quantum resources cannot be established, we can still analyze the energy output based the structure of the initial state.
Consider the initial mixed state $\rho_{AB}$ with the “X”-structure, as expressed in the form
$\rho_{AB}^{X}=\begin{pmatrix}a&0&0&\alpha e^{i\beta}\\ 0&b&\delta e^{i\epsilon}&0\\ 0&\delta e^{-i\epsilon}&c&0\\ \alpha e^{-i\beta}&0&0&d\end{pmatrix}$, (13)
where all parameters are real and satisfy the normalization condition $a+b+c+d=1$ [58, 59]. Suppose that we adopt the Hamiltonian of system with the form in Eq. (1). Initially, prior to the QET protocol, the system in the state $\rho_{AB}^{X}$ has the energy
$E_{0}\left(\rho_{AB}^{X}\right)=(a+b-c-d)\varepsilon_{A}+(a-b+c-d)\varepsilon_{B}\\ +4\kappa(\delta\cos\epsilon+\alpha\cos\beta).$ (14)
The energy of the system after the measurements $P_{A}(u)$ is given by
$E_{A}\left(\rho_{AB}^{X}\right)=(a-b+c-d)\varepsilon_{B}+4\kappa(\delta\cos\epsilon+\alpha\cos\beta),$ (15)
while the injected energy is $E_{A}-E_{0}=-(a+b-c-d)\varepsilon_{A}$. Finally, the energy of the system after Bob’s correction $U_{B}(u)$ is
$E_{B}\left(\rho_{AB}^{X}\right)\\ =\left((a-b+c-d)\varepsilon_{B}+4\kappa(\delta\cos\epsilon+\alpha\cos\beta)\right)\cos 2\theta\\ -2\left((-a+b-c+d)\kappa+\varepsilon_{B}(\delta\cos\epsilon+\alpha\cos\beta)\right)\sin 2\theta.$ (16)
The energy output $E_{\text{out}}=E_{A}-E_{B}$ is given by
$E_{\text{out}}\left(\rho_{AB}^{X}\right)=D\sin 2\theta-F(1-\cos 2\theta),$ (17)
where
$D=2(-a+b-c+d)\kappa+\varepsilon_{B}(\delta\cos\epsilon+\alpha\cos\beta),$
$F=-(a-b+c-d)\varepsilon_{B}-4\kappa(\delta\cos\epsilon+\alpha\cos\beta).$
It is evident that the output of energy is dependent on the parameter $\theta$. The maximal value is given by
$\tan(2\theta_{1})=\frac{D}{F}\quad\text{or}\quad\tan\left(2\theta_{2}+\frac{\pi}{2}\right)=\frac{D}{F},$ (18)
and the corresponding energy output is
$E^{\text{max}}_{\text{out}}\left(\rho_{AB}^{X}\right)=\sqrt{D^{2}+F^{2}}-F.$ (19)
Note that the optimal $\theta$, giving the maximal energy output, is not unique, but the maximum value of $E_{\text{out}}$ remains the same.
II.3 Environments and Bloch-Redfield equation
The mixed initial state $\rho_{AB}$ arises due to environmental influence. We consider a scenario where each qubit couples to a separate environment, potentially with distinct temperatures or chemical potentials. This configuration is designed to ensure the system device remains reusable rather than disposable. The environment not only represents an unavoidable factor but also serves to reset the apparatus. Specifically, we employ the steady state of the system. After protocol completion, the environment resets the system, enabling the next operational cycle.
The total Hamiltonian combining the system and the environment is given by
$H=H_{AB}+H_{R}+H_{I},$ (20)
where $H_{AB}$ is the Hamiltonian of the two interacting qubits, as defined in Eq. (1). The free Hamiltonian of the reservoirs, $H_{R}$, is
$H_{R}=\sum_{k_{A}}\omega_{k_{A}}b_{k_{A}}^{\dagger}b_{k_{A}}+\sum_{k_{B}}\omega_{k_{B}}b_{k_{B}}^{\dagger}b_{k_{B}},$ (21)
where $b_{k_{A}}$ ($b_{k_{A}}^{\dagger}$) and $b_{k_{B}}$ ($b_{k_{B}}^{\dagger}$) are the annihilation (creation) operators for the $k$-th mode with frequencies $\omega_{k_{A}}$ and $\omega_{k_{B}}$ of the reservoirs coupled to qubits $A$ and $B$, respectively. The qubit-reservoir interaction under the rotating wave approximation is
$H_{I}=\sum_{k_{A}}g_{k_{A}}\left(\sigma^{-}_{A}b_{k_{A}}^{\dagger}+\sigma^{+}_{A}b_{k_{A}}\right)\\ +\sum_{k_{B}}g_{k_{B}}\left(\sigma^{-}_{B}b_{k_{B}}^{\dagger}+\sigma^{+}_{B}b_{k_{B}}\right),$ (22)
where $g_{k_{A}}$ and $g_{k_{B}}$ are qubit-reservoir coupling strengths. In the eigenbasis of $H_{S}$ (1), interaction Hamiltonian $H_{I}$ can be rewritten as
$H_{I}=\sum_{k_{A}}g_{k_{A}}(\eta_{A}+\xi_{A})b_{k_{A}}^{\dagger}\\ +\sum_{k_{B}}g_{k_{B}}(\eta_{B}+\xi_{B})b_{k_{B}}^{\dagger}+\text{H.c.},$ (23)
where $\eta_{A,B},\xi_{A,B}$ are transition operators given by
$\eta_{A}$ $=\sin(\phi_{1}+\phi_{2})(|E_{3}\rangle\langle E_{4}|-|E_{1}\rangle\langle E_{2}|),$
$\eta_{B}$ $=\cos(\phi_{1}-\phi_{2})(|E_{3}\rangle\langle E_{4}|+|E_{1}\rangle\langle E_{2}|),$
$\xi_{A}$ $=\cos(\phi_{1}+\phi_{2})(|E_{2}\rangle\langle E_{4}|+|E_{1}\rangle\langle E_{3}|),$
$\xi_{B}$ $=\sin(\phi_{1}-\phi_{2})(|E_{2}\rangle\langle E_{4}|-|E_{1}\rangle\langle E_{3}|).$ (24)
The corresponding transition frequencies are
$\varepsilon_{\pm}=\sqrt{\Omega+4\kappa^{2}}\pm\sqrt{\Delta+4\kappa^{2}}.$ (25)
Here $\varepsilon_{-}$ corresponds to the transitions from the state $|E_{2}\rangle$ to $|E_{1}\rangle$ and the state $|E_{4}\rangle$ to $|E_{3}\rangle$. The transition frequency $\varepsilon_{+}$ corresponds to the transitions from the state $|E_{4}\rangle$ to $|E_{2}\rangle$ and the state $|E_{3}\rangle$ to $|E_{1}\rangle$.
The Born-Markov quantum master equation in the interaction picture is given by [40, 41]
$\frac{d\tilde{\rho}_{AB}}{dt}=-\int_{0}^{\infty}ds~\mathrm{Tr}_{R}\left[\tilde{H}_{I}(t),[\tilde{H}_{I}(t-s),\tilde{\rho}_{AB}\otimes\tilde{\rho}_{R}]\right],$ (26)
[…]
$\frac{d\rho_{AB}}{dt}=-i[H_{AB},\rho_{AB}]+\sum_{j=A,B}\mathcal{D}_{j}(\rho_{AB}),$ (27)
where $\mathcal{D}_{j}(\rho_{AB})$ is the dissipator given by
$\mathcal{D}_{j}(\rho_{AB})=\\ \alpha_{j}(\varepsilon_{-})(\eta_{j}^{\dagger}\rho_{AB}\eta_{j}+\eta_{j}^{\dagger}\rho_{AB}\xi_{j}-\eta_{j}\eta_{j}^{\dagger}\rho_{AB}-\xi_{j}\eta_{j}^{\dagger}\rho_{AB})\\ +\alpha_{j}(\varepsilon_{+})(\xi_{j}^{\dagger}\rho_{AB}\xi_{j}+\eta_{j}^{\dagger}\rho_{AB}\xi_{j}-\xi_{j}\xi_{j}^{\dagger}\rho_{AB}-\eta_{j}\xi_{j}^{\dagger}\rho_{AB})\\ +\beta_{j}(\varepsilon_{-})(\eta_{j}\rho_{AB}\eta_{j}^{\dagger}+\eta_{j}\rho_{AB}\xi_{j}^{\dagger}-\eta_{j}^{\dagger}\eta_{j}\rho_{AB}-\xi_{j}^{\dagger}\eta_{j}\rho_{AB})\\ +\beta_{j}(\varepsilon_{+})(\xi_{j}\rho_{AB}\xi_{j}^{\dagger}+\eta_{j}\rho_{AB}\xi_{j}^{\dagger}-\xi_{j}^{\dagger}\xi_{j}\rho_{AB}-\eta_{j}^{\dagger}\xi_{j}\rho_{AB})\\ +\text{H.c.}$ (28)
Here the coefficients $\alpha_{j}(\varepsilon)$ and $\beta_{j}(\varepsilon)$ are the dissipation rates, given by
$\alpha_{j}(\varepsilon)=$ $\gamma_{j}(\varepsilon)n_{j}(\varepsilon),$
$\beta_{j}(\varepsilon)=$ $\gamma_{j}(\varepsilon)(1\pm n_{j}(\varepsilon)),$ (29)
where the coupling spectrum $\gamma_{j}(\varepsilon)$ is
$\gamma_{j}(\varepsilon)=\pi\sum_{k_{j}}|g_{k_{j}}|^{2}\delta(\varepsilon-\omega_{k_{j}})$ (30)
and $n_{j}(\varepsilon)$ is the Bose-Einstein (minus sign) or the Fermi-Dirac (plus sign) distribution
$n_{j}(\varepsilon)=\frac{1}{e^{(\varepsilon-\mu_{j})/T_{j}}\mp 1}.$ (31)
The sign of $\beta_{j}(\varepsilon)$ is plus for the bosonic reservoirs, while it is minus for fermionic reservoirs. Parameters $T_{j}$ and $\mu_{j}$ are the equilibrium temperatures and chemical potentials of $j$-th reservoir, respectively. For bosonic reservoirs, such as photon or phonon baths, the particle number is not conserved with a vanishing chemical potential. Since the Bloch-Redfield equation is based on the assumption that the interaction between the system and the environment is weak, we can further assume that the coupling spectra with different frequencies are much less than the energy scale of the two qubits, namely $g_{k}\ll\varepsilon_{A},\varepsilon_{B}$. Therefore, it is reasonable to view $g_{k}$ as constants (independent of the transition frequencies $\varepsilon_{\pm}$).
The two-qubit steady state can be solved by reformulating the Bloch-Redfield equation in the Liouville space [28, 60]. It corresponds to the eigenstates of the superoperator with the zero eigenvalue. Because the four eigenstates of $H_{AB}$ is in $X$ form (13), the steady state either in the Hamiltonian eigenbasis or the local basis is an $X$ density operator.
III Quantum energy teleportation under equilibrium environments
We separately discuss the influence of bosonic and fermionic equilibrium environments on QET in Secs. III.1 and III.2 respectively.
III.1 Equilibrium bosonic environments
Figure 2: Energy output of steady states in the equilibrium bosonic reservoirs and the corresponding population of $|E_{1}\rangle$. ($\text{a}_{1}$) Energy output when the energy levels are set as $\varepsilon=0.1~(\text{black solid line})$, 1(red dashed line)1~($\text${red dashed line}), 5(green dot line)5~($\text${green dot line}) and 10(blue dashed dot line)10~($\text${blue dashed dot line}). ($\text{a}_{2}$) The population of $|E_{1}\rangle$ corresponding to ($\text{a}_{1}$). ($\text{b}_{1}$) Energy output when the temperatures are set as T=0.1(black solid line)T=0.1~($\text${black solid line}), 1(red dashed line)1~($\text${red dashed line}), 5(green dot line)5~($\text${green dot line}) and 10(blue dashed dot line)10~($\text${blue dashed dot line}). ($\text{b}_{2}$) The population of $|E_{1}\rangle$ corresponding to ($\text{b}_{2}$). The other parameters are set as $\kappa=1$ and $g_{A}=g_{B}=0.05$.
Suppose the initial state is a mixed state of eigenstates of $H_{AB}$, such as a thermal state. The energy output is less than the weighted sum of the maximum $E_{\text{out}}$ from each eigenstate. This suggests that a more efficient QET may exist for mixed states. Recall that we have analytical results for the energy output of the four distinct eigenstates given by Eq. (12). Therefore, the environmental influence on energy output can be analyzed by studying how the weighting factors of these eigenstates vary.
The system population in bosonic reservoirs is not reversible; the ground state has the maximum population, and in the high-temperature limit, all states exhibit equal population. This results in a relatively minor contribution of the excited states to energy output. Although the maximum $E_{\text{out}}$ of excited states when considered in isolation is greater than that of the ground state, this advantage is insufficient to counterbalance the negative contributions from the ground state at the same parameter $\theta$ when taking the populations. Consequently, the $E_{\text{out}}$ predominantly depends on the ground state.
In the bosonic case, if the parameters of both reservoirs are identical, the system reaches the steady state of $H_{AB}$. We denote the equilibrium temperature as $T=T_{1}=T_{2}$. The energy output exhibits a brief plateau as temperature increases, followed by a quick decline, as shown in Fig. 2 $(\text{a}_{1})$. The plateau duration increases for systems with higher energy levels $\varepsilon$, which can be explained by the suppressed thermal excitation at low temperatures, as illustrated in Fig. 2 $(\text{a}_{2})$.
Furthermore, the energy output does not scale linearly with the energy level $\varepsilon$, as shown in Fig. 2 $(\text{b}_{1})$. When $\varepsilon$ increases at a fixed equilibrium temperature, the population of $|E_{1}\rangle$ approaches unity, as shown in Fig. 2 $(\text{b}_{2})$. The energy output initially grows with $\varepsilon$ but declines once the population of state $|E_{1}\rangle$ saturates. When the population of $|E_{1}\rangle$ approaches 1, the energy output $E_{\text{out}}(|E_{1}\rangle)$ has simple form as $4\kappa^{2}/\sqrt{4\varepsilon^{2}+4\kappa^{2}}$ (obtained form Eq. 19). Therefore, as $\varepsilon$ increases, $E_{\text{out}}$ will diminish.
III.2 Equilibrium fermionic environments
Figure 3: (a) Energy output of steady states with increasing chemical potential in fermionic reservoirs. The parameters are set as $\varepsilon=1$ and $\theta=\theta_{1}$ (black solid line), $\varepsilon=1$ and $\theta=\theta_{2}$ (red dashed line), $\varepsilon=3$ and $\theta=\theta_{1}$ (green solid line), $\varepsilon=3$ and $\theta=\theta_{2}$ (blue dashed line). (b) The population of state $|E_{4}\rangle$ with $\mu$. The parameters are set as $\varepsilon=1$ (black solid line) or $\varepsilon=3$ (red dashed line). The other parameters are set as $\kappa=1$, $T=T_{A}=T_{B}=1$, and $g_{A}=g_{B}=0.05$.
Consider fermionic reservoirs with identical equilibrium temperatures and chemical potentials $\mu=\mu_{A}=\mu_{B}$. When the equilibrium chemical potential surpasses the system’s energy levels, the population of the highest excited state $|E_{4}\rangle$ dominates over the other three states. In this regime, the QET protocol with the parameter $\theta=\theta_{1}$ given by Eq. (18) yields negative energy output. Conversely, using the optimal parameter $\theta=\theta_{2}$ (Eq. 18) significantly enhances the energy output while strictly maintaining energy conservation (see Fig. 3).
We examine the energy output variation for two distinct parameters $\theta$ as a function of the equilibrium chemical potential $\mu$, as depicted in Fig. 3 (a). Increasing the chemical potential induces population inversion among the eigenstates. At sufficiently high chemical potentials, the excited states acquire significant populations. Figure 3 (b) shows that the population of the highest excited state $|E_{4}\rangle$ grows monotonically with $\mu$. When $\mu$ surpasses the system’s energy levels, the energy output curves for the $\theta_{1}$ and $\theta_{2}$ protocols undergo an abrupt interchange. This indicates that as the state $|E_{4}\rangle$ population dominates, the mixed-state energy output behavior converges toward that of the pure state $E_{\text{out}}(|E_{4}\rangle)$.
IV Quantum energy teleportation under nonequilibrium environments
We separately discuss the influence of bosonic and fermionic nonequilibrium environments on QET in Secs. IV.1 and IV.2 respectively.
IV.1 Nonequilibrium bosonic environments
Figure 4: Energy output of steady states under the nonequilibrium bosonic environments or the energy detuning, and the corresponding population of eigenstates. ($\text{a}_{1}$) The average temperatures are set as $\bar{T}=0.5$ (black solid line), $\bar{T}=2$ (red dashed line), and $\bar{T}=5$ (green dot line). The energy levels are set as $\varepsilon_{A}=\varepsilon_{B}=2$. ($\text{a}_{2}$) The population of state $|E_{1}\rangle$ (in solid line) and state $|E_{2}\rangle$ (in dashed line) corresponding to ($\text{a}_{1}$). ($\text{b}_{1}$) The energy levels are set as $\bar{\varepsilon}=1$ with $T=0.5$ (black solid line), $\bar{\varepsilon}=2$ with $T=0.5$ (red dashed line), and $\bar{\varepsilon}=2$ with $T=2$ (green dot line). ($\text{b}_{2}$) The population of state $|E_{1}\rangle$ (in solid line) and state $|E_{2}\rangle$ (in dashed line) corresponding to ($\text{b}_{1}$). The other parameters are set as $\kappa=1$ and $g_{A}=g_{B}=0.05$.
When the temperatures of two reservoirs are not the same, we have a nonequilibrium environment. We denote the temperature difference $\Delta T=T_{A}-T_{B}$ to quantify the nonequilibriumness. Under nonequilibrium bosonic environments, the energy output $E_{\text{out}}$ decreases as the average temperature $\bar{T}=(T_{A}+T_{B})/2$ increases, as demonstrated in Fig. 4 $(\text{a}_{1})$. At lower average temperatures, the energy output is reduced with increasing temperature difference $|\Delta T|$, as illustrated by the curve for $T=0.5$ in Fig. 4 $(\text{a}_{1})$. Correspondingly, the population of the ground state remains above 0.9, significantly exceeding that of other excited states, as shown in Fig. 4 $(\text{a}_{2})$ (black solid line). As $|\Delta T|$ increases, the population of the ground state decreases, mirroring the trend observed in $E_{\text{out}}$.
When $\bar{T}=2$, the population of the ground state remains above 0.65, while the population of the first excited state remains above 0.2. Unlike the population of the ground state, the population of the first excited state increases as $|\Delta T|$ rises. Consequently, the influence of the first excited state on $E_{\text{out}}$ becomes more pronounced, leading to a scenario where $E_{\text{out}}$ increases with $|\Delta T|$. When $\Delta T$ is relatively extreme, the decrease in the population of the ground state does not correspond to the increase in the population of the first excited state. Higher excited states also occupy a portion, which results in a reduction of the corresponding $E_{\text{out}}$.
As the average temperature continues to rise to $5$, the two qubits exhibit no entanglement, and $E_{\text{out}}$ approaches 0 with minimal variations. Overall, the temperature difference significantly reduces $E_{\text{out}}$ at low temperatures, while at higher temperatures, it exerts a slight enhancement on $E_{\text{out}}$.
To further enhance the nonequilibrium phenomenal, we consider the QET protocol with the initial state from two detuned qubits, namely $\Delta\varepsilon\neq 0$ with $\Delta\varepsilon=\varepsilon_{A}-\varepsilon_{B}$, while the average is set as $\bar{\varepsilon}=(\varepsilon_{A}+\varepsilon_{B})/2$. As illustrated in Fig. 4 $(\text{b}_{1})$, the energy output $E_{\text{out}}$ increases with an increase in the detuning $\Delta\varepsilon$. As $\Delta\varepsilon$ increases, the energy level of qubit A rises, allowing for a greater amount of energy to be injected. Conversely, Bob’s energy level imposes a limit on the maximum energy output. Therefore, when $\Delta\varepsilon$ is small, $E_{\text{out}}$ increases with the detuning $\Delta\varepsilon$. However, once $E_{B}$ (energy after Bob preforms the correction $U_{B}(u)$) decreases below a certain threshold, the energy output decreases as $\Delta\varepsilon$ increase. The population of state $|E_{1}\rangle$ decreases with $|\Delta\varepsilon|$ while the population of state $|E_{2}\rangle$ increases as shown in Fig. 4 $(\text{b}_{2})$. The influence of temperature on $E_{\text{out}}$ remains significant. The increase in temperature enhances the excitations within the system while reducing the energy output.
The combination of detuning within the system and the nonequilibrium environments results in a significant enhancement of $E_{\text{out}}$ (with fixed average energy level $\bar{\varepsilon}$ and fixed average temperatures $\bar{T}$). Specifically, as qubit $A$ has a higher energy level and couples to higher temperature reservoirs, the energy output can be greatly enhanced, as shown Fig. 5. The influence of the nonequilibrium environments on the detuning two qubits is asymmetrical. In the region where $E_{\text{out}}$ is enhanced, the population of the ground state also increases, since the qubit with higher energy level coupled to the higher temperature reservoir, makes it more difficult for the system to become excited.
Figure 5: Energy output of steady states of two detuned qubits under the bosonic nonequilibrium environments. The parameters are set as $\bar{\varepsilon}=(\varepsilon_{A}+\varepsilon_{B})/2=2$, $\bar{T}=(T_{A}+T_{B})/2=0.5$, $\kappa=1$, and $g_{A}=g_{B}=0.05$.
IV.2 Nonequilibrium fermionic environments
Figure 6: The energy output of steady states under the fermionic reservoirs with nonequilibrium temperatures, and the corresponding population of eigenstates. The chemical potentials are set as (a) $\mu=1$, (b) $\mu=2$, and (c) $\mu=8$. ($\text{a}_{2}$) The population of state $|E_{1}\rangle$ (in black solid line) and state $|E_{2}\rangle$ (in red dashed line) corresponding to ($\text{a}_{1}$). ($\text{b}_{2}$) The population of state $|E_{2}\rangle$ (in black solid line) and state $|E_{4}\rangle$ (in red dashed line) corresponding to the ($\text{b}_{1}$). ($\text{c}_{2}$) The population of state $|E_{2}\rangle$ (in black solid line) and state $|E_{4}\rangle$ (in red dashed line) corresponding to the ($\text{c}_{1}$). The other parameters are set as $\kappa=1$, $\bar{T}=1$, and $g_{A}=g_{B}=0.05$.
When the two qubits coupled with nonequilibrium fermionic environments, we separately discuss the influence of the temperature difference and the chemical potential difference on the energy output. When the chemical potential is relatively low, the energy output from the system is predominantly determined by the population of state $|E_{1}\rangle$ and the $E_{\text{out}}$ is low, as shown in Fig. 6 ($\text{a}_{1}$). The temperature difference $\Delta T$ can reduce the excited states population and enhance the population of state $|E_{1}\rangle$, leading to an increase of $E_{\text{out}}$, as shown in Fig. 6 ($\text{a}_{2}$). But in the extreme case of $|\Delta T|$, the population of $|E_{1}\rangle$ decreases, which corresponds to the reduction of $E_{\text{out}}$, as shown in Fig. 6 ($\text{a}_{1}$).
When the chemical potential is comparatively high, the energy output is governed by the population of state $|E_{4}\rangle$, correspondingly $E_{\text{out}}$ is significantly enhanced, as illustrated in Fig. 6 ($\text{c}_{1}$). The increase of temperature difference $\Delta T$ increases $E_{\text{out}}$ by facilitating transitions from the state $|E_{2}\rangle$ to $|E_{4}\rangle$. For the extreme nonequilibrium cases (with large $|\Delta T|$), the population of state $|E_{4}\rangle$ decreases, which corresponds to the reduction of $E_{\text{out}}$, as shown in Fig. 6 ($\text{c}_{2}$).
In cases of moderate chemical potential (comparable to the energy level), the energy output has two distinct scenarios. The parameter $\theta$ of QET protocol (in the correction operator $U_{B}(u)$) changes from $\theta_{2}$ to $\theta_{1}$ with the increasing $|\Delta T|$, as depicted in Fig. 6 ($\text{b}_{1}$). In this configuration, the thermal effect and the particle exchange from the nonequilibrium fermionic environments have a combined excitation effect. In other words, the population of state $|E_{4}\rangle$ is sufficiently large, corresponding to the QET protocol with the parameter $\theta_{2}$. The temperature difference leads to a reduction in the population of state $|E_{4}\rangle$ due to the decreased excitation level of the qubit at the low-temperature reservoir, resulting in a decrease in $E_{\text{out}}$, as shown in Fig. 6 ($\text{b}_{2}$). Furthermore, the reduction in population due to the presence of chemical potential is transferred to state $|E_{2}\rangle$ rather than the ground state.
When the chemical potentials of two reservoirs are not same, $\Delta\mu=\mu_{A}-\mu_{B}$, it is essential to analyze the situation on a case-by-case basis (dependent on the value of average chemical potential). When the average chemical potential $\bar{\mu}=(\mu_{A}+\mu_{B})/2$ is relatively low, the primary effect of $|\Delta\mu|$ is to push the system from the ground state to the first excited state, as shown in Fig. 7 ($\text{a}_{2}$). In this context, the energy output $E_{\text{out}}$ mainly depends on the population of the ground state. Therefore, as $|\Delta\mu|$ increases, the energy output subsequently decreases as demonstrated in Fig. 7 ($\text{a}_{1}$).
When the average chemical potential rises, the population of the highest excited state gradually increases. Energy output $E_{\text{out}}$ is enhanced when $|\Delta\mu|$ is small; while reduced when $|\Delta\mu|$ is far away from the equilibrium, as shown in Fig. 7 ($\text{b}_{1}$). The chemical potential difference $|\Delta\mu|$ can enhance the population of state $|E_{4}\rangle|$ when it is less than approximately $0.5$, as shown in Fig. 7 ($\text{b}_{2}$). As $|\Delta\mu|$ increases, the qubit connected to the lower chemical potential reservoir tends to be de-excited, reflected on the population of state $|E_{2}\rangle|$. When the chemical potential difference is large, the population of all eigenstates becomes equal, resulting in mutual cancellation of energy extraction between the different states, leading to a vanishing $E_{\text{out}}$.
In cases of high average chemical potential, as shown in Fig. 7 ($\text{c}_{1}$), the variation in $E_{\text{out}}$ relies on the population of the state $|E_{4}\rangle|$. Since in high chemical potential case the population of state $|E_{4}\rangle|$ is predominant, an increase in the chemical potential difference results in a smaller population differences, as shown in Fig. 7 ($\text{c}_{2}$), thereby leading to a decrease in energy output. The chemical potential difference can only decrease the energy output.
Figure 7: The energy output of steady states under the fermionic reservoirs with nonequilibrium chemical potential, and the corresponding population of eigenstates. The average chemical potentials are set as (a) $\bar{\mu}=1$, (b) $\bar{\mu}=6$, and (c) $\bar{\mu}=8$. ($\text{a}_{2}$) The population of state $|E_{1}\rangle$ (in black solid line) and state $|E_{2}\rangle$ (in red dashed line) corresponding to ($\text{a}_{1}$). ($\text{b}_{2}$) The population of state $|E_{1}\rangle$ (in black solid line), state $|E_{2}\rangle$ (in red dashed line), state $|E_{3}\rangle$ (in green dot line), and state $|E_{2}\rangle$ (in blue dash-dot line) corresponding to ($\text{b}_{1}$). ($\text{c}_{2}$) The population of state $|E_{1}\rangle$ (in black solid line), state $|E_{2}\rangle$ (in red dashed line), state $|E_{3}\rangle$ (in green dot line), and state $|E_{2}\rangle$ (in blue dash-dot line) corresponding to ($\text{c}_{1}$). The other parameters are set as $\kappa=1$, $T_{A}=T_{B}=1$ and $g_{A}=g_{B}=0.05$.
If we consider the detuned two qubits with nonzero $\Delta\varepsilon=\varepsilon_{A}-\varepsilon_{B}$, the energy output exhibits asymmetry respect to $\Delta\varepsilon$. When the chemical potential is relatively low, the energy difference $\Delta\varepsilon$ primarily affects the population of states $|E_{1}\rangle$ and $|E_{2}\rangle$, as shown in Fig. 8 ($\text{a}_{1}$). When $\Delta\varepsilon$ is non-zero, the qubit with lower energy level becomes more easily excited, resulting in an increase in the population of the first excited state $|E_{1}\rangle$. When $\Delta\varepsilon$ is small, the population of state $|E_{1}\rangle$ is predominant, and the energy output initially increases with $\Delta\varepsilon$ before subsequently decreasing as shown in Fig. 8 ($\text{a}_{2}$). The initial increase of $E_{\text{out}}$ is due to the enhancement of $\varepsilon_{A}$, while the decrease correlates with a reduction in the population of $|E_{1}\rangle$. Conversely, when $\Delta\varepsilon$ is large, the population of $|E_{2}\rangle$ becomes dominant. At this point, the behavior of $E_{\text{out}}$ closely follows that of $E_{\text{out}}(|E_{2}\rangle)$, being enhanced by $\Delta\varepsilon$ due to the increasing population of $|E_{2}\rangle$.
When the chemical potential is relatively high, the two-qubit system is predominantly in excited states, as illustrated in Fig. 8 ($\text{b}_{1}$). For large detuning $\Delta\varepsilon$, the individual qubit becomes difficult for excitation. In such instances, the population distribution primarily concentrates in states $|E_{2}\rangle$ and $|E_{4}\rangle$, as illustrated in Fig. 8 ($\text{b}_{2}$). As $|\Delta\varepsilon|$ increases, the population of state $|E_{4}\rangle$ compensates for that of state $|E_{2}\rangle$. When the population of state $|E_{2}\rangle$ becomes dominant, the QET protocol switchs the parameter from $\theta_{1}$ to $\theta_{2}$, and the curve for $E_{\text{out}}$ experiences a sudden change.
Figure 8: Energy output of steady states with detuned energy levels. The parameters are set as (a) $\mu=1$ and (b) $\mu=8$. ($\text{a}_{2}$) The population of state $|E_{1}\rangle$ (in black solid line) and state $|E_{2}\rangle$ (in red dashed line) corresponding to ($\text{a}_{1}$). ($\text{b}_{2}$) The population of state $|E_{2}\rangle$ (in black solid line) and state $|E_{4}\rangle$ (in red dashed line) corresponding to ($\text{b}_{1}$). The other parameters are set as $\kappa=1$, $\bar{\varepsilon}=2$, $T_{A}=T_{B}=1$, and $g_{A}=g_{B}=0.05$.
Figure 9: Energy output of steady states with two nonequilibrium parameters. The parameters are set as ($\text{a}_{1}$): $\bar{T}=0.5,\bar{\mu}=1~\text{and}~\varepsilon_{A}=\varepsilon_{B}=2$; ($\text{a}_{2}$): $\bar{T}=0.5,\bar{\mu}=8~\text{and}~\varepsilon_{A}=\varepsilon_{B}=2$; $b_{1}$: $\bar{\varepsilon}=2,\bar{T}=0.5~\text{and}~\mu_{A}=\mu_{B}=1$; $b_{2}$: $\bar{\varepsilon}=2,\bar{T}=0.5~\text{and}~\mu_{A}=\mu_{B}=8$; $c_{1}$: $\bar{\varepsilon}=2,\bar{\mu}=1~\text{and}~T_{A}=T_{B}=0.5$ and $c_{2}$: $\bar{\varepsilon}=2,\bar{\mu}=6~\text{and}~T_{A}=T_{B}=0.5$.The other parameters are set as $\kappa=1$, and $g_{A}=g_{B}=0.05$.
When both the temperatures and chemical potentials are nonequilibrium, for small average chemical potentials, changes in $E_{\text{out}}$ correlates with variations in the population of the ground state. The combination of a high (low) temperature reservoir with a low (high) chemical potential leads to an increase in the population of the ground state, thereby enhancing the energy output as shown in Fig. 9 ($\text{a}_{1}$). In scenarios characterized by high average chemical potential, the population of $E_{\text{out}}$ aligns with the population of state $|E_{4}\rangle$. In this case, the influence of the temperature difference is small, and $E_{\text{out}}$ is primarily governed by the chemical potential difference and the population of $E_{\text{out}}$ resembles that of the scenario in which only the chemical potential is out of equilibrium, as shown in Fig. 9 ($\text{a}_{2}$). Overall, $E_{\text{out}}$ has an enhancement in specific nonequilibrium regions.
When the detuned two qubits coupled to the nonequilibrium environments, the distribution of $E_{\text{out}}$ is no longer centrally symmetric around the equilibrium position. When the temperatures of two reservoirs are different, in the case of low chemical potential, as the energy level difference $\Delta E$ increases, $E_{\text{out}}$ becomes significantly elevated when $\Delta T$ is considerably large due to the increase in the population of the first excited state, as shown in Fig. 9 ($\text{b}_{1}$). Additionally, the high energy level qubit coupled to the reservoir with a high temperature, can further enhance the population of state $|E_{2}\rangle$. In scenarios with relatively high chemical potential, the magnitude of $E_{\text{out}}$ is primarily determined by the population of the highest excited state. Consequently, $E_{\text{out}}$ can be notably amplified in the upper left region of Fig. 9 ($\text{b}_{2}$).
When the detuned two qubits coupled to the nonequilibrium environments with different chemical potentials, maintaining Bob’s qubit at a high chemical potential is advantageous for energy output.
In scenarios with low average chemical potential, the distribution of $E_{\text{out}}$ clearly represents the combination of the effects of energy detuning and the chemical potential difference, as demonstrated in Fig.9 ($\text{c}_{1}$). Conversely, at the high average chemical potential, the qubit with a high (low) energy level coupled a high (low) chemical potentials gives a larger population of state $|E_{4}\rangle$, which in turn can enhance the energy output, as shown in Fig.9 ($\text{c}_{2}$). A larger $\varepsilon_{B}$ has the potential to extract more energy, thereby leading to greater energy output in the lower left quadrant of Fig.9 ($\text{c}_{2}$).
V Conclusions
In our study, we explore the impact of both equilibrium and nonequilibrium parameters on the QET, based on a two-qubit model coupled with two separate environments. By discussing the energy output behaviors, we qualitatively analyze how equilibrium and nonequilibrium environments influence the energy output by affecting the population of energy eigenstates within a mixed state.
In the bosonic reservoirs, we find that when the energy level of qubit A is higher and in contact with a higher temperature reservoir, it leads to an increase in input energy, thereby enhancing the energy output. Nonequilibrium conditions primarily influence the energy output by affecting the population of the ground state. By analyzing the combined effects of temperature differences and detuned energy levels on the system, the energy output can be significantly enhanced in regions $\Delta T>0$ with $\Delta\varepsilon>0$, compared to the equilibrium cases.
For fermionic reservoirs, the scenario is more complex. At a low chemical potential, the energy output is major determined by the ground state population, which is similar to the behavior observed in bosonic reservoirs. However, in the cases of high chemical potential, the population of the highest excited state is large, which is the main factor affecting energy output.
For fermionic reservoirs, temperature differences can generally enhance the energy output, while the chemical potential difference mainly reduces it. We have also considered the scenarios of combining the nonequilibrium temperatures and the chemical potentials, as well as two detuned qubits coupled with the nonequilibrium environments with a temperature or chemical potential difference. The energy output $E_{\text{out}}$ can be enhanced in certain parameter regions that are far from the equilibrium position. Overall, the nonequilibrium conditions for both bosonic reservoirs and fermionic reservoirs can improve the performance of QET.
Note that the QET protocols for the four different eigenstates have different optimal control operations, indicating that the protocol can not fully extract energy from mixture of eigenstates. It implies that there may be more effective QET strategies for mixed states. In ideal situation, the new energy extraction protocol could be applicable to all eigenstates and enable greater energy retrieval from mixed states. If achieved, this would also provide valuable insights into the quantum resources upon which QET depends.
Acknowledgements.
X. K. Yan thanks NSF 12234019 for support. K. Zhang is supported by the National Natural Science Foundation of China under Grant Nos. 12305028 and 12247103, China Postdoctoral Science Foundation under Grant Number 2025M773421, Scientific Research Program Funded by Education Department of Shaanxi Provincial Government (Program No.24JP186), and the Youth Innovation Team of Shaanxi Universities.